OpenStudy (anonymous):

Simplify the complex fraction. @kx2bay

OpenStudy (anonymous):

\[\frac{ \frac{ \frac{ 5x^8 }{ y^4 } }{ 10x^7 } }{ y^5 }\]

hartnn (hartnn):